OpenStudy (anonymous):

Which expression is equivalent to sec (-theta )? Would this be 1/-cos(theta)?

OpenStudy (anonymous):

no that is incorrect

OpenStudy (anonymous):

\[\sec(-\theta)=\frac{1}{\cos(-\theta)}=\frac{1}{\cos(\theta)}\] because cosine is even so \[\cos(-x)=\cos(x)\][

OpenStudy (anonymous):

ohhhhh I see.

OpenStudy (anonymous):

Mason owns a collection of hardback books in a series about American history. Each book has dimensions as shown below. There are 7 books in the series, stacked side by side. What is the volume of the entire collection? a) one hundred sixty-one cubic inches b) eight hundred forty cubic inches c) nine hundred twenty-four cubic inches d) nine hundred sixty cubic inches

OpenStudy (anonymous):

help plzzz illl give u a medal